OpenStudy (mpari13):

for what values of a and b is the following function continuous at every x? F(x)= -4 x <= -3 ax-b -3=1

OpenStudy (anonymous):

this is much easier than it looks replace \(x\) by \(-3\) and set the result equal to \(-4\) then replace \(x\)by \(1\) and set the result equal to \(2\) solve for \(a\) and \(b\)

OpenStudy (mpari13):

so your saying set a(-3)-b = -4?

OpenStudy (mpari13):

ohh okay you mean set the x in f(x)= -4 and then =2

OpenStudy (anonymous):

\[-3a-b=-4\\ a-b=2\]

OpenStudy (anonymous):

i.e. \[a\times (-3)-b=-4\\ a(1)-b=2\]

OpenStudy (anonymous):

then solve the system \[-3a-b=-4\\ a-b=2\] for \(a\) and \(b\)